Hello, tutors! Please help me with the problem down below :)

The time t (in seconds) it takes a dropped object to fall h feet is given by t=1/4√h.

a.) Use a graphing calculator to graph the function. Describe the domain and range.

b.) It takes about 13.17 seconds for a ball dropped from the golden gate bridge in san francisco to reach water below. about how high is the bridge from the water?

Answer by Alan3354(69443)

The time t (in seconds) it takes a dropped object to fall h feet is given by t=1/4√h.

a.) Use a graphing calculator to graph the function. Describe the domain and range.
---------
ok, I did that.
Do you have one?
====================
b.) It takes about 13.17 seconds for a ball dropped from the golden gate bridge in san francisco to reach water below. about how high is the bridge from the water?
---
Is it t = (1/4)sqrt(h) ?
Or t = 1/(4sqrt(h)) ?